OpenStudy (anonymous):

function in R F(x)=1/2x^2 find set points That can create 2 tangent to the curve C(f) orthogonal

OpenStudy (anonymous):

first u need to kow what is orthoganal, do u have any idea?

OpenStudy (anonymous):

y1=a1x+b1 y2=a2x+b2 if orthogonal a1*a2=-1

OpenStudy (anonymous):

okay from that i understand that a1 and a2 are our gradients right? so if a1*a2=-1 the two tagent are perpendicular to each other

OpenStudy (anonymous):

ok but how ??

OpenStudy (anonymous):

sorry it is perpendicular not orthogonal .... i understand english a little

OpenStudy (anonymous):

 i can't draw but ur graph is something like that, both sides of the graph are identic we are asked to find the two points let call the A and B

OpenStudy (anonymous):

yes it is thx very much

OpenStudy (anonymous):

which language do u speak better?

OpenStudy (anonymous):

arabic

OpenStudy (anonymous):

oh i speak french so we can only stick to english, back to our question after delopping, i let f(x)=y.... so y=1/2x^2 yx^2=1/2

OpenStudy (anonymous):

ah ok je suis algerienne je parle francais ok ok pas probleme ...... merci

OpenStudy (anonymous):

peux tu trouver la solution a ton probleme maintenant maintenant ou on doit continuer?

OpenStudy (anonymous):

ooooook merci

OpenStudy (anonymous):

que ce tu as trouver?